# Eclipsed X-ray Bursts from Magnetar SGR J1935+2154 and the Fireball Measurements

### v1

@@ 2.2 Eclipse Modeling
$\vec{e}_{\mathrm{burst}}$ $=$ $\left(\sin\chi\cos\xi-\cos\chi\sin\xi\cos\gamma,\sin\xi\sin\gamma,\cos\chi\cos\xi+\sin\chi\sin\xi\cos\gamma\right),$ (2)
where $\chi$ is the angle between the observer and the spin axis and $\xi$ is the angle between the spin axis and the fireball. Since the fireball co-rotates with the magnetar, the vector $\vec{e}_{\mathrm{burst}}$ rotates around $\vec{e}_{\mathrm{spin}}$. Therefore, the rotational phase is given by $\gamma\left(t\right)=\Omega t+\phi$, with $\Omega=2\pi/P$ and $\phi$ being related to the choice of zero-time during each burst.
@@
For the selected bursts, we assume their emitting flux to be constant except for the fast rising and decay phases. Specifically, we write the intrinsic light curve as follows:
$F_{\mathrm{int}}=F_{p}\times\Theta\left(t-t_{1}\right)\Theta\left(t_{2}-t\right),$ (7)
where $t_{1}$ and $t_{2}$ are the starting and ending times of the burst plateau, $F_{p}$ is the peak fluxes of each burst, and $\Theta(t)$ is the step function. Therefore, the observed flux is simply given by
@@
We fit the light curves of the four selected bursts, assuming an intrinsically steady fireball, as shown by the dashed lines in Figure 4, with the corresponding posterior distributions shown in Figure 5. The corresponding parameter values with errors are listed in Table 1. In calculations, the intrinsic radii of the fireballs are taken beforehand according to the spectral fittings. The viewing angles derived from the four eclipsed bursts are found to be all concentrated around a common value  [^1] The common value is estimated from the mean of individual posteriors and includes uncertainties, $\bar{\chi}=\frac{1}{4}\sum_{i=B}^{E}\bar{\chi}_{i},\quad\sigma^{2}=\frac{1}{4}\sum_{i=B}^{E}\left[\sigma^{2}_{i}+(\bar{\chi}_{i}-\bar{\chi})^{2}\right],$ (9) where $\bar{\chi}_{i=\mathrm{B,...E}}$ and $\sigma_{i}^{2}$ is the posterior mean and variance of $i$-th burst, respectively. of $\bar{\chi}=17^{\circ}\pm 10^{\circ}$, as shown in Figure 6, indicating the self-consistency of our modeling. It is further established that the pulsar (neutron star) viewing angle can be determined through a novel eclipse-based method, in contrast to the conventional rotating vector model which relies on radio polarization measurements. The results further show that $d\gg R_{\mathrm{ns}}$ and $d\gg l$, which indicate the emitted fireballs of these bursts are far away from but do not adhere to the magnetar’s surface. This is consistent with the implication of the CRSFs. Finally, the distance $d$ is much smaller than the radius of the light cylinder as $R_{\rm LC}=cP/2\pi=1.5\times 10^{5}$ km, suggesting that the fireballs are located deep in the magnetosphere.
Note that, in our model, we do not take into account the gravitational lensing effect for simplicity, which can somewhat increase the brightness of a distant light source if the light can graze near the surface of the magnetar. However, given that the fireball is comparable to or several times larger than the magnetar in size, we should treat the fireball as an extended source rather than a point. In such a case, since different differential elements on the source have different distances to the caustic, the magnification of the total brightness of the fireball would be averaged over the entire extended surface. As a result, the brightness variation arising from the distance change of the fireball to the caustic would be much smaller than that observed for a point source (S. Dall’Osso et al., 2024).
@@
Table 1: The observational data utilized and the eclipse model fitting parameters. The errors are given at the $1\sigma$ confidence level. The symbols $l$, $d$, $\chi$, and $\xi$ are defined as illustrated in Figure 1. $\phi$ is the phase corresponding to the starting time, $R_{\mathrm{ns}}$ is the radius of the magnetar, and $F_{\mathrm{p}}$ is the peak flux of the bursts.
@@
3 Discussion and Summary
The distance measurements of the fireballs to the magnetar surface provide a very stringent constraint on the trigger mechanism of the bursts. Generally speaking, highly twisted magnetic fields in the magnetosphere of magnetars are very likely to extend from the stellar interior, where the toroidal component could dominate. In other words, the energy release of magnetars is ultimately a result of the diffusion of magnetic helicity from the interior to the exterior. The connection between the internal and external fields exerts vast stress on the solid crust of magnetar. Then, along with the spin-down of magnetar, the shear stress is gradually accumulated to finally exceed the yield limit of the crust, leading to plastic deformation of the crust and triggering a starquake. In principle, the released strain energy can heat the deforming crustal region, leading to the formation of a hot spot and even a fireball trapped by local magnetic field loops near the surface. However, such a picture is now disfavored for these eclipsed plateau bursts, as their distances are about five times larger than the stellar radius. Instead, the fireballs are inferred to be suspended in the magnetosphere, suggesting them to originate from local magnetic reconnection of some magnetic ropes/knots in the magnetosphere. The magnetic ropes/knots may form as the cumulative product of long-term crustal deformation. Thus, the occurrence of such catastrophic magnetic reconnection is not necessarily linked to an individual starquake, even though the energy released ultimately derives from the cumulative effect of numerous previous starquakes.
In more detail, we can envision the following physical scenario. As a magnetic reconnection occurs in the magnetosphere, an energy of about $E_{\rm iso}\sim 10^{41}$ erg can be released suddenly. Then, a great number of electron and positron pairs can be excited in the reconnection zone and flow into a larger scale region, forming of a freely and adiabatically expanded fireball. The number density of the fireball can be given by (L. D. Landau & E. M. Lifshitz, 1980; C. Thompson & R. C. Duncan, 1995)
@@ 3 Discussion and Summary
where the internal temperature of the fireball $T_{\rm in}$ can be roughly estimated to be $k_{\rm B}T_{\rm in}=k_{\rm B}({3E_{\rm iso}/4\pi l^{3}a})^{1/4}=68(E_{\rm iso}/10^{41}{\rm erg})^{1/4}(l/20{\rm km})^{-3/4}\rm keV$ with $a$ being the radiation constant. The fireball expansion could be quickly suppressed if the pressure of the surrounding large-scale magnetic fields can finally exceed the thermal pressure of the fireball, which requires $B\gtrsim 1.6\times 10^{11}\>({l}/{20\>\mathrm{km}})^{-3/2}({E_{\mathrm{iso}}}/{10^{41}\>\mathrm{erg}})^{1/2}\>\mathrm{G}$ (C. Thompson & R. C. Duncan, 1995). By using a simple dipolar field configuration, the field strength at the distance $d$ can be estimated to $B(d)\sim 3.0\times 10^{12}({d}/{50\>\mathrm{km}})^{-3}\mathrm{G}$, which is obviously strong enough to stop the fireball expansion, where a reference stellar radius $R_{\mathrm{ns}}=12$ km is adopted and the surface field strength is taken as $B_{\mathrm{p}}=2.2\times 10^{14}\>\mathrm{G}$ (G. L. Israel et al., 2016). Therefore, the fireball can finally be trapped by the magnetic fields and suspended in the magnetosphere for a period.
@@
The rationality of a suspended fireball may also be understood from the perspective of interpreting the key properties of the CRSF absorption, including its central energy, line width, and absorption depth. (i) With the magnetic field strength at a reference distance of 50 km, we can determine the position of the first harmonic of CRSFs in the spectra of the bursts as: $E_{\mathrm{cyc}}=\hbar{eB/m_{\rm e}c}=35.3\left({d/50\ {\rm km}}\right)^{-3}\rm keV$, where $e$ and $m_{\rm e}$ are respectively the charge and mass of electrons, $c$ is the speed of light, and the used formula is valid for the case of $B\ll B_{\rm c}=4.4\times 10^{13}$ G with $B_{\rm c}$ being the critical value for Landau quantization. Although very simple, this estimate of $E_{\mathrm{cyc}}$ can be very consistent with the values obtained from the spectral fittings, further indicating the self-consistency of the model. However, strictly speaking, some inconsistencies still exit, since the fireball distances inferred from the light curve fittings (see Table 1) are actually somewhat different from the reference value. Here, the crux of the problem may lie in the complication of the magnetic field configuration of the magnetar, which distinguishes it from the traditional dipolar scenario and complicates the determination of both $B(d)$ and $E_{\rm cyc}$ at a given distance. (ii) It is expected that the widths of CRSF is at least larger than that determined by the Doppler effect as (P. Meszaros & W. Nagel, 1985; A. K. Harding & J. K. Daugherty, 1991; O. Nishimura, 2003; G. Schönherr et al., 2007): $\sigma=({k_{\rm B}T/m_{\rm e}c^{2}})^{1/2}E_{\rm cyc}|\cos\theta|=4.9{\rm keV}\left({k_{\rm B}T/{10\rm keV}}\right)^{1/2}\left({E_{\rm cyc}/35\rm keV}\right)|\cos\theta|$, where $\theta$ is the angle between the photon direction and the magnetic field vector. This value is smaller than, but of the same order of magnitude as, the CRSF widths derived from the spectral fittings. However, since the fireball is actually distributed within a spatial range of a few tens of kilometers, the magnetic field strength in the vicinity of the fireball should at least vary on an amplitude of $(d+l)/(d-l)]^{3}$. Given this significant variation in the field strength, it would be not surprising to observe a CRSF line a few times broader than the Doppler broadening. (iii) In the outer region adjacent to the fireball surface where the CRSF could be formed, the column density of electrons and positrons can be estimated by $\mu n_{\rm GJ}\Delta$, where $\mu$ is the multiplicity, $n_{\rm GJ}=\Omega B/(2\pi ce)$ is the Goldreich-Julian density, and $\Delta$ is the thickness of this line-formation region. Then, following Reference J. C. L. Wang et al. (1993), the optical depth of the CRSF absorption can be expressed as $\tau_{\mathrm{CRSF}}=0.3\left({\mu/10^{3}}\right)\left({\Delta/5\rm km}\right)\left({k_{\rm B}T/\rm 10keV}\right)^{-1/2}$, which could provide a basic explanation for the optical depths derived from the spectral fittings.
@@
Finally, the dynamical evolution of a fireball is determined by the competition between the internal thermal pressure and external magnetic pressure. Supposing that a burst is triggered impulsively, a plateau emission would occur during a pressure balance. Then, as the majority of the internal energy of the fireball is lost, the balance would be disrupted, and thus the plateau emission ceases. Therefore, we can express the duration of the plateau emission in terms of the diffusion timescale of thermal photons in the fireball as
$t_{\rm diff}\sim{l^{2}\over\lambda}=3.3\times 10^{4}\left({k_{\rm B}T_{\rm in}\over 1\rm keV}\right)^{3/2}\exp\left(-{511{\rm keV}\over k_{\rm B}T_{\rm in}}\right)~{\rm s},$ (11)
@@
where $\lambda=1/(\sigma_{\rm T}n_{\pm})$ is the mean free path of the thermal photons with $\sigma_{\rm T}$ being the Thomson cross section and $l=20$ km is adopted for giving the numerical coefficient. However, this timescale is very sensitive to the determination of the internal temperature of the fireball because of the presence of the exponential term. Specifically, by solving equation $t_{\rm diff}\sim 1$ s, we can obtain $k_{\rm B}T_{\rm in}=32.7$ keV for $l=20$ km, which is lower than that inferred from the average internal energy density. Here, it should be noted that the actual temperature within the fireball must exhibit a decreasing distribution from the inside out, which is determined by thermal transfer within the fireball. Given this, it is not surprising to obtain different temperatures from different estimation methods, not to mention that these approaches are rather crude. Then, undoubtedly, a precise calculation of the burst duration requires a detailed characterization of the thermal diffusion. In addition, more complexity can arise from the unclear trigger mechanism of the bursts. Given the complication of locally twisted magnetic fields, magnetic reconnection is likely to occur repeatedly and with complicated dynamical behavior rather than impulsively as supposed above. In this case, the pressure balance of the fireball cannot keep steady and, thus, an evolving and even multi-peaked light curve rather than a plateau would be obtained. Such complicated bursts can be easily found from the longest bursts of SGR J1935+2154, with a larger number than plateau bursts. Since we cannot know the dynamical evolution of the fireballs of these complicated bursts, it is actually impossible to identify eclipse candidates from them, even though their spectral properties are not obviously different from the plateau ones. Furthermore, it is indicated that plateau bursts are actually very rare, which seems sound because of their strict requirement on the burst trigger mechanism. In any case, these facts make it difficult for us to verify the eclipse model from the perspective of number statistics.

### v2

@@ II.2 Eclipse Modeling
$\vec{e}_{\mathrm{burst}}$ $=$ $\left(\sin\chi\cos\xi+\cos\chi\sin\xi\cos\gamma,\sin\xi\sin\gamma,\cos\chi\cos\xi-\sin\chi\sin\xi\cos\gamma\right),$ (2)
where $\chi$ is the angle between the observer and the spin axis and $\xi$ is the angle between the spin axis and the fireball. Since the fireball co-rotates with the magnetar, the vector $\vec{e}_{\mathrm{burst}}$ rotates around $\vec{e}_{\mathrm{spin}}$. Therefore, the rotational phase is given by $\gamma\left(t\right)=\Omega t+\phi$, with $\Omega=2\pi/P$ and $\phi$ being the azimuth angle of the fireball, which is related to the choice of starting-time during each burst.
@@
For the selected bursts, we assume their emission flux to be constant except during the fast-rising and decay phases. Specifically, we write the intrinsic light curve of the burst as follows:
$F_{\mathrm{int}}=F_{\mathrm{p}}\times\Theta\left(t-t_{1}\right)\Theta\left(t_{2}-t\right),$ (7)
where $t_{1}$ and $t_{2}$ are the starting and ending times of the burst plateau, $F_{\mathrm{p}}$ is the peak flux of the fireball emission without eclipse, and $\Theta(t)$ is the step function. Therefore, the observed flux is simply given by
@@
With this model, we fit the light curves of the four eclipsed bursts, assuming an intrinsically steady fireball, as shown by the dashed lines in Figure 4, with the corresponding posterior distributions shown in Figure 5. The corresponding parameter values with errors are listed in Table 1. In the calculations, the intrinsic radii of the fireballs are taken from the spectral fits. The viewing angles derived from the four eclipsed bursts are found to be all concentrated around a common value  [^2] The common value is estimated from the mean of individual posteriors and includes uncertainties, $\bar{\chi}=\frac{1}{4}\sum_{i=B}^{E}\bar{\chi}_{i},\quad\sigma^{2}=\frac{1}{4}\sum_{i=B}^{E}\left[\sigma^{2}_{i}+(\bar{\chi}_{i}-\bar{\chi})^{2}\right],$ (9) where $\bar{\chi}_{i=\mathrm{B,...E}}$ and $\sigma_{i}^{2}$ is the posterior mean and variance of $i$-th burst, respectively. of $\bar{\chi}=17^{\circ}\pm 10^{\circ}$, as shown in Figure 6, indicating the self-consistency of our modeling. We stress that the pulsar (neutron star) viewing angle can be determined through this novel eclipse-based method, in contrast to the conventional rotating vector model that relies on radio polarization measurements. The derived values of the parameter $\xi$ are relatively large, because only for those fireballs located at the opposite side of the observer that would be eclipsed by the magnetar itself. Our results further show that $d\gg R_{\mathrm{ns}}$ and $d\gg l$, which indicate the emitted fireballs of these bursts are far away from but do not adhere to the magnetar’s surface. This is consistent with the implication of the CRSFs. Finally, the distance $d$ is much smaller than the radius of the light cylinder as $R_{\rm LC}=cP/2\pi=1.5\times 10^{5}$ km, suggesting that the fireballs are located deep in the magnetosphere.
It is necessary to note that in our co-rotating spherical fireball model, we focus on the intermediate plateau phase of the burst, during which the emitting region approximately co-rotates with the magnetosphere. Physically, the fireball is expected to undergo rapid expansion and subsequent cooling and contraction, as in classical trapped fireball scenarios, but these processes correspond to the short rise and decay phases and occur on timescales much shorter than the plateau duration. We therefore approximate the fireball as a quasi-steady, uniform emitter during this phase, which allows us to investigate how global magnetospheric rotation and geometrical viewing angles imprint on the observed flux modulation without modeling the detailed microphysical evolution. Furthermore, we do not account for gravitational lensing in our model, which can slightly increase the brightness of a distant light source if the light grazes the magnetar’s surface. However, given that the fireball is comparable to or several times larger than the magnetar in size, we should treat the fireball as an extended source rather than a point source. In such a case, since different differential elements on the source have different distances to the caustic, the magnification of the total brightness of the fireball should be averaged over the entire extended surface. As a result, the brightness variation arising from the distance change of the fireball to the caustic would be much smaller than that for a point source (Dall’Osso et al., 2024).
@@
Table 1: The observational data utilized and the eclipse model fitting parameters. The errors are given at the $1\sigma$ confidence level. The symbols $l$, $d$, $\chi$, and $\xi$ are defined as illustrated in Figure 1. $\phi$ is the phase corresponding to the starting time, $R_{\mathrm{ns}}$ is the radius of the magnetar, and $F_{\mathrm{p}}$ is the peak flux of the fireball emission without eclipse.
@@
III Discussion
The distance measurements of the fireballs to the magnetar surface provide a very stringent constraint on the trigger mechanism of the bursts. Generally speaking, highly twisted magnetic fields in the magnetosphere of magnetars are very likely to extend from the stellar interior, where the toroidal component could dominate. In other words, the energy release of magnetars is ultimately a result of the diffusion of magnetic helicity from the interior to the exterior. The connection between the internal and external fields exerts vast stress on the solid crust of magnetar. Then, along with the spin-down of magnetar, the shear stress is gradually accumulated to finally exceed the yield limit of the crust, leading to plastic deformation of the crust and triggering a starquake. In principle, the released strain energy can heat the deforming crustal region, leading to the formation of a hot spot and even a fireball trapped by local magnetic field loops near the surface.
However, such a picture is now disfavored for these eclipsed plateau bursts, as their distances are approximately five times the stellar radius. Instead, the fireballs are inferred to be suspended in the magnetosphere, suggesting that they originate from local magnetic reconnection of magnetic ropes or knots. The magnetic ropes/knots may form as the cumulative product of long-term crustal deformation. Thus, the occurrence of such catastrophic magnetic reconnection is not necessarily linked to an individual starquake, even though the energy released ultimately derives from the cumulative effect of numerous previous starquakes. It is worth pointing out that our conclusions are only applicable to the selected bursts with intrinsic plateau emission and observed eclipsing feature. So far, our analysis does not rule out the starquake scenario that also forms a fireball.
Furthermore, we can envision the following physical scenario. As a magnetic reconnection occurs in the magnetosphere, an energy of about $E_{\rm iso}\sim 10^{41}$ erg can be released suddenly. Then, a great number of electron and positron pairs can be excited in the reconnection zone and flow into a larger-scale region, forming a magnetically trapped fireball. The number density of the fireball can be given by (Landau and Lifshitz, 1980; Thompson and Duncan, 1995)
@@ III Discussion
where $e$ and $m_{\rm e}$ are respectively the charge and mass of electrons, $c$ is the speed of light. The internal temperature of the fireball $T_{\rm in}$ can be roughly estimated to be $k_{\rm B}T_{\rm in}=k_{\rm B}({3E_{\rm iso}/4\pi l^{3}a})^{1/4}=68(E_{\rm iso}/10^{41}{\rm erg})^{1/4}(l/20{\rm km})^{-3/4}\rm keV$ with $a$ being the radiation constant. The fireball expansion could be quickly suppressed if the pressure of the surrounding large-scale magnetic fields can finally exceed the thermal pressure of the fireball, which requires $B\gtrsim 1.6\times 10^{11}\>({l}/{20\>\mathrm{km}})^{-3/2}({E_{\mathrm{iso}}}/{10^{41}\>\mathrm{erg}})^{1/2}\>\mathrm{G}$ (Thompson and Duncan, 1995). By using a simple dipolar field configuration, the field strength at the distance $d$ can be estimated to $B(d)\sim 3.0\times 10^{12}({d}/{50\>\mathrm{km}})^{-3}\mathrm{G}$, which is obviously strong enough to stop the fireball expansion, where a reference stellar radius $R_{\mathrm{ns}}=12$ km is adopted and the surface field strength is taken as $B_{\mathrm{p}}=2.2\times 10^{14}\>\mathrm{G}$ (Israel et al., 2016). Therefore, the fireball can finally be trapped by the magnetic fields and suspended in the magnetosphere for a time interval.
@@
The rationality of a suspended fireball may also be understood from the perspective of interpreting the key properties of the CRSF absorption, including its central energy, line width, and absorption depth: (i) With the magnetic field strength at a reference distance of 50 km, we can determine the position of the first harmonic of CRSFs in the spectra of the bursts as: $E_{\mathrm{cyc}}=\hbar{eB/m_{\rm e}c}=35.3\left({d/50\ {\rm km}}\right)^{-3}\rm keV$, which is valid for the case of $B\ll B_{\rm c}=4.4\times 10^{13}$ G with $B_{\rm c}$ being the critical value for Landau quantization. Although very simple, this estimate of $E_{\mathrm{cyc}}$ can be very consistent with the values obtained from the spectral fittings, further indicating the self-consistency of the model. However, strictly speaking, some inconsistencies still exist, since the fireball distances inferred from the light curve fittings (see Table 1) are actually somewhat different from the reference value. Here, the crux of the problem may lie in the complication of the magnetic field configuration of the magnetar, which distinguishes itself from the traditional dipolar scenario and complicates the determination of both $B(d)$ and $E_{\rm cyc}$ at a given distance. (ii) It is expected that the widths of CRSF are at least larger than that determined by the Doppler effect as (Meszaros and Nagel, 1985; Harding and Daugherty, 1991; Nishimura, 2003; Schönherr et al., 2007): $\sigma=({k_{\rm B}T/m_{\rm e}c^{2}})^{1/2}E_{\rm cyc}|\cos\theta|=4.9{\rm keV}\left({k_{\rm B}T/{10\rm keV}}\right)^{1/2}\left({E_{\rm cyc}/35\rm keV}\right)|\cos\theta|$, where $\theta$ is the angle between the photon direction and the magnetic field vector. This value is smaller than, but of the same order of magnitude as, the CRSF widths derived from the spectral fittings. However, since the fireball is actually distributed within a spatial range of a few tens of kilometers, the magnetic field strength in the vicinity of the fireball should at least vary on an amplitude of $(d+l)/(d-l)]^{3}$. Given this significant variation in the field strength, it would be not surprising to observe a CRSF line a few times broader than the Doppler broadening. (iii) In the outer region adjacent to the fireball surface where the CRSF could be formed, the column density of electrons and positrons can be estimated by $\mu n_{\rm GJ}\Delta$, where $\mu$ is the multiplicity, $n_{\rm GJ}=\Omega B/(2\pi ce)$ is the Goldreich-Julian density, and $\Delta$ is the thickness of this line-formation region. Then, following Wang et al. (1993), the optical depth of the CRSF absorption can be expressed as $\tau_{\mathrm{CRSF}}=0.3\left({\mu/10^{3}}\right)\left({\Delta/5~\rm km}\right)\left({k_{\rm B}T/\rm 10~keV}\right)^{-1/2}$, which could provide a basic explanation for the optical depths derived from the spectral fittings.
@@
The dynamical evolution of a fireball is determined by the competition between the internal thermal pressure and external magnetic pressure. Supposing that a burst is triggered impulsively, a plateau emission would occur during a pressure balance. Then, as the majority of the internal energy of the fireball is lost, the balance would be disrupted, and thus the plateau emission ceases. Therefore, we can express the duration of the plateau emission in terms of the diffusion timescale of thermal photons in the fireball as
$t_{\rm diff}\sim{l^{2}\over c\lambda}=3.3\times 10^{4}\left({k_{\rm B}T_{\rm in}\over 1~\rm keV}\right)^{3/2}\exp\left(-{511{\rm keV}\over k_{\rm B}T_{\rm in}}\right)~{\rm s},$ (11)
@@
where $\lambda=1/(\sigma_{\rm T}n_{\pm})$ is the mean free path of the thermal photons with $\sigma_{\rm T}$ being the Thomson cross section and $l=20$ km is adopted for giving the numerical coefficient. However, this timescale is very sensitive to the determination of the internal temperature of the fireball because of the presence of the exponential term. Specifically, by solving equation $t_{\rm diff}\sim 1$ s, we can obtain $k_{\rm B}T_{\rm in}=32.7$ keV for $l=20$ km, which is lower than that inferred from the average internal energy density. Here, it should be noted that the actual temperature within the fireball must exhibit a decreasing distribution from the inside out, which is determined by thermal transfer within the fireball. Given this, it is not surprising to obtain different temperatures from different estimation methods, not to mention that these approaches are rather crude. Then, undoubtedly, a precise calculation of the burst duration requires a detailed characterization of the thermal diffusion.
In addition, more complexity can arise from the unclear trigger mechanism of the bursts. Given the complication of locally twisted magnetic fields, magnetic reconnection is likely to occur repeatedly and with complicated dynamical behavior rather than impulsively as supposed above. In this case, the pressure balance of the fireball cannot keep steady and, thus, an evolving and even multi-peaked light curve rather than a plateau would be obtained. Such complicated bursts can be easily found from the longest bursts of SGR J1935+2154, with a larger number than plateau bursts. Since we cannot know the dynamical evolution of the fireballs of these complicated bursts, it is actually impossible to identify eclipse candidates from them, even though their spectral properties are not obviously different from the plateau ones.
Furthermore, it is indicated that plateau bursts are actually very rare, which seems sound because of their strict requirement on the burst trigger mechanism. In any case, these facts make it difficult for us to verify the eclipse model from the perspective of number statistics.
IV Summary
In this paper, we identified one representative plateau-like burst among the long bursts of SGR J1935+2154 detected by GECAM and Fermi/GBM, and found four candidate bursts showing eclipse-like features. We established the co-rotating fireball geometric eclipse model and fitted the light curves of four eclipse bursts. We derived the viewing angle related to the spin axis with $\chi\approx 17^{\circ}$, and the fireball is located at several NS radii away from the magnetar. This distance indicates that the fireball is suspended in the magnetosphere, where the magnetic field strength is estimated to be $B\sim 3\times 10^{12}~\mathrm{G}$. Moreover, we find CRSF features in the spectra of X-ray bursts, which also implies a magnetic field strength similar to that inferred from the light-curve distances. These results not only provide smoking-gun evidence of the co-rotating fireball, which accounts for the energetic X-ray emission during the magnetar activity, but also support an origin of the fireball from magnetic reconnection in the magnetosphere rather than directly from a crustal starquake. We also note that this new eclipse-based method provides a novel tool to constrain the viewing angle and could be applied to other magnetar bursts.
This work is supported by the National Key R&D Program of China (2021YFA0718500), the National SKA Program of China (2020SKA0120300), the National Natural Science Foundation of China (grant Nos. 12393811, 12393812, and 12273042), the Strategic Priority Research Program, the Chinese Academy of Sciences (Grant No. XDB0550300) and the China’s Space Origins Exploration Program. The GECAM (HuaiRou-1) mission is supported by the Strategic Priority Research Program on Space Science of the Chinese Academy of Sciences. We appreciate the public data and software of Fermi/GBM.